# 1-2-3 Check: Enhancing Contextual Privacy in LLM via Multi-Agent Reasoning

### v1

@@ 5.1 Multi-Agent and Information Flow
To systematically evaluate how different agent configurations balance privacy preservation and completeness of public content, we conducted experiments on single-agent, two-agent, and three-agent pipelines with different information flow configurations. We show our information flow exploration results in table˜1, with detailed description and case studies are in appendix˜B. We additionally include a Chain-of-Thought single-agent baseline, which follows the same instructions as the three-agent setup, which detecting events, classifying them as private or public, and generating meeting summaries.
@@
Our results is as table˜1 overall, the multi-agent framework balances secret leakage with public information omission, enhancing both data security and retention. Across both backbones, multi-agent pipelines substantially reduce our composite error (Leaks Secret or Omits Info↓) relative to single-agent baselines. For LLaMA-3.1-70B-Instruct, the best configuration is a three-agent, the information flow that the checker only delivers public information (20.5), improving over the single-agent baseline by 28.0-28.0. For GPT-4o, the best configuration is a three-agent, Annotate-Private, no transcript flow ($17.5$), improving over the single-agent baseline ($26.3$) by 8.8-$8.$ Notably, moving from two to three agents yields additional gains for both models (LLaMA: 24.5 $\rightarrow$ 20.5; GPT-4o: 21.5 $\rightarrow$ 17.5 depending on flow), indicating that introducing a dedicated checker and clearly separating agent duties improves the privacy utility trade-off provided that each role is well-defined and properly aligned with the task structure.
@@
For LLaMA-3.1-70B-Instruct, providing the executor with the meeting transcript leads to a clear advantage for the Public Only strategy: both in the two-agent (Composite error: $24.5$) and three-agent (Composite: $20.5$) settings, this approach outperforms Annotate Private ($34.5$ and $42.0$, respectively). Here, restricting the checker to only forward public information, combined with transcript access, allows the executor to recover useful details and reduces the risk of private information leakage. Importantly, the Public Only strategy also simplifies the executor’s task by eliminating the need for the model to further distinguish between private and public content during its own processing. However, in this setting, if the meeting transcript is withheld from the executor, the performance of Public Only degrades sharply, where omissions of public information increase substantially (e.g., three-agent omission: $19.5$ with transcript vs. $26.0$ without), and composite errors rise (three-agent: 20.5 $\rightarrow$ 32.0). This suggests that LLaMA relies heavily on the transcript to supplement the limited information from the checker, and without it, relevant content is easily lost.
For GPT-4o, the interplay is more balanced. With transcript access, Annotate Private achieves the lowest leakage (5.0), while Public Only achieves the lowest omission (9.0); their composite errors are comparable (24.5 vs. 23.5). When the transcript is withheld, however, the three-agent Annotate Private setting yields the best overall performance (Composite: 17.5), with much lower omission (10.0) than Public Only without transcript (8.5), and still reasonable control of leakage (8.5 vs. 15.0). This indicates that the stronger GPT-4o executor can effectively use checker tags to reconstruct public content, even without seeing the original transcript, and is less prone to privacy leaks than LLaMA in the same setup.
@@ 5.2 Three Agent Information Propagation
To further understand the conditions under which the three-agent framework improves privacy preservation, we conducted some quantitative analysis of information flow across agents in different settings.
@@
Our results (see fig.˜15) demonstrate that the Assistant agent alone is insufficient for reliably identifying public information, necessitating a dedicated Checker agent for effective public information restoration. In this respect, GPT-4o outperforms Llama3.1-70B (55.0% vs. 46.5%) in recovering public content, and the Executor also contributes to the restoration process.
The Checker agent highlights the pronounced superiority of GPT-4o in privacy preservation compared to Llama3.1-70B, as evidenced by a substantially lower privacy leakage rate (33.0% vs. 79.5%). Moreover, GPT-4o achieves a much lower overall information flow leakage rate across the entire pipeline (20.5% vs. 67.0%), which directly substantiates our claim and elucidates why Llama3.1-70B, when used as a base model without providing the meeting transcript to the Executor, suffers significant performance degradation. In contrast, GPT-4o’s performance remains stable,or even improves, under these constraints, illustrating the effectiveness of the Checker’s filtering mechanism and reducing dependence on the downstream agent’s ability to reconstruct prior context (see fig.˜16).
Overall, the composite quality score further reinforces this pattern: while both models exhibit comparable performance at the Assistant stage, Llama3.1-70B’s quality declines substantially at the Checker stage, whereas GPT-4o not only maintains but improves its performance, with a further increase observed at the Executor stage.
@@
To further dissect the role of checker annotation strategies, we conducted a detailed information propagation analysis, comparing Privacy Annotated and Public Only settings across agent stages.
For Llama-3.1-70B-Instruct, our results see(fig.˜17) reveal a striking difference: at the checker stage, the proportion of private information included in the Public Only pipeline is substantially higher than in the Privacy Annotated configuration (83.5% vs. 31.0%). This suggests that without explicit privacy annotations, the Public Only setting often fails to sufficiently filter sensitive content. For public information, as expected, the Public Only setting consistently yields higher mention rates across all agent stages compared to the Privacy Annotated setting.
In contrast, GPT-4o exhibits an even more pronounced effect (see fig.˜18): when the checker operates in Privacy Annotated mode, the rate of private information mentions is much higher than in Public Only (83% vs. 25.5%). Similarly, for public information, the Public Only setting again achieves higher mention rates than Privacy Annotated.
These results highlight the importance of using a strong model as the checker for effective privacy filtering, while also suggesting that multi-agent pipelines should adaptively assign stronger models to critical validation tasks and reserve weaker models for less sensitive stages, with additional verification steps as needed.
@@ 5.3 Three Agent In Privacylens
Figure˜2 presents the privacy–utility Pareto frontier across all evaluated configurations. We observe a consistent pattern: multi-agent information flows substantially increase privacy preservation with only minor reductions in helpfulness. For both GPT-4.1 and GPT-4o, moving from Single agent to Privacy Annotation or Public Only achieves PPR gains of 15–19 percentage points while maintaining high AHS (from 2.56 to 2.53). Notably, the Privacy Annotation variant offers the best overall Pareto-optimal balance, achieving both high utility and robust privacy protection.
@@ 6 Conclusion
We introduced a multi-agent approach that partitions the tasks of event extraction, classification, and final summary generation among separate agents, addressing the limitations of a single LLM operating alone. Experimental results and ablation studies confirm that our multi-agent pipelines significantly reduce private information leakage without substantially compromising the quality. Notably, we observed performance differences across various information flow configurations and uncovered the multi-agent system’s capacity for self-correction. Our findings highlight the importance of modular, intermediate validation steps in complex, context-dependent scenarios.
@@ 7 Limitation
Our experiments focus on meeting summarization and privacy scenarios defined by the ConfAIde (Mireshghallah et al., 2023) and PrivacyLens (Shao et al., 2025) benchmark. Although our multi-agent architecture demonstrates strong performance in these controlled benchmarks, domain adaptation remains non-trivial. The Extractor’s event schema and the Checker’s privacy rules must be handcrafted for each application domain—medical, legal, or financial—each of which has distinct notions of sensitive versus permissible disclosures. This tailoring demands specialized prompt engineering, rule design, and potential collaboration with domain experts. To date, no publicly available benchmarks systematically evaluate multi-agent or information-flow architectures for privacy in domains such as healthcare, finance, or legal document processing. Future work should prioritize building such benchmarks, defining domain-specific privacy norms which would help to better showing the generalizability of our multi-agent system.
Residual Hallucination and Inference.
Although the framework mitigates direct leakage by controlling event flow between agents, LLMs can still infer private details from partial context or generate hallucinations that indirectly breaches privacy. For instance, even if the transcript redacts certain details, a language model might infer or reconstruct them from other cues. Our current approach focuses on explicit event classification but does not robustly account for inference-based leaks in more complex scenarios or highly entangled data.
Acknowledge
This material is based upon work supported by the Defense Advanced Research Projects Agency (DARPA) under Agreement No. HR00112490410, and was also partly funded by Fujitsu.
@@ A.1 Prompt For ConfAIde
Instruction: You are an AI assistant that helps summarize conversations from a meeting based on the transcript. The summary will be automatically sent to all the attendees of the meeting, no matter when they joined.
@@ A.3 Hyperparameter
To investigate information propagation, we adopted the following decoding configurations across all experiments for both LLaMA-3-Instruct and GPT-4 models. We set the decoding temperature to 1.0 and used a top-p value of 1.0 to enable unconstrained sampling from the full probability distribution. To avoid introducing biases into the generation process, we did not apply frequency or presence penalties (both set to 0.0). For reproducibility, a fixed random seed of 99 was used throughout. Additionally, we generated 10 samples per input prompt to account for generation variability. When evaluating the generalizability of our three-agent framework, we aligned all model decoding temperatures to 1.0 to ensure a fair comparison with the thinking models (i.e., o3 and o4-mini) which use the same sampling configuration.
@@ B.1 Case Studies of the Single Agent Baseline Framework
We present a case study on baseline-agent results in table˜4. In this scenario, we observe that the single agent fails to identify private information, leading to the private information leakying. (specifically, the private event: a surprise party).
@@ B.2 Case Studies of the Two-Agent Framework
We present a case study on two-agent information flow in table˜5. In this scenario, we observe that the assistant agent fails to identify private information, leading to the propagation of erroneous data (specifically, the private event: a surprise party) to the subsequent executor agent.
@@ B.3 Case Studies of the Three-Agent Framework
Qualitative case studies, such as the example in table˜6 and table˜7, demonstrate the effectiveness of our three-agent configuration. We observe that the Checker Agent, whether filtering events to include only public information or annotating private information, consistently achieves strong performance in identifying and classifying events. This careful segmentation of sensitive versus non-sensitive content results in an Executor Agent output that reliably maintains privacy while conveying essential public details.
@@
Based on the results from the private-only and public-only configurations, we observed that the Executor can autonomously refine the information transmitted by the Checker Agent by leveraging cues from the meeting transcript during the summary generation process. Consequently, we explored a scenario in which the Executor is provided solely with the Checker’s information—omitting the meeting transcript entirely—to assess both the Executor’s performance in this context and the accuracy of the Checker’s public-only setup. Our qualitative investigation and case study (see table˜8) shows that the Checker often fails to accurately identify all public information, which increases the likelihood of public details being omitted in the final output.
Appendix C Chain of Thought Single Agent
Appendix D Three-Agent Performance Analysis
@@
To probe the single-agent’s upper bound, in ConfAIde, we introduce a more targeted prompt—Enhanced Single Agent (see appendix˜A)—designed to maximize public-information inclusion. fig.˜13 illustrates the trend across four configurations: Baseline Single Agent, Enhanced Single Agent, Three-Agent Privacy Annotation, and Three-Agent Public-Only. Across all models, our Three-Agent framework consistently outperforms the Baseline Single Agent. When compared against the Enhanced Single Agent prompt, the Three-Agent framework delivers especially pronounced gains for open-source models, where it matches or slightly exceeds the performance of the closed-source thinking models. Beyond public inclusion, the Three-Agent framework also shows strong privacy protection across most backbones. Its overall performance is comparable to that of the best closed-source thinking models, while significantly outperforming single-agent settings on non-thinking models.
@@
The privacy-annotated information and public-only information flow’s results of PrivacyLens are shown in fig.˜14. We observe a clear trade-off between helpfulness and privacy preservation in our three-agent framework. Our quantitative results show that the binary helpfulness rate and the average helpfulness score remain virtually unchanged when moving from a single-agent to a three-agent configuration across all six models. In stark contrast, privacy preservation metrics improve dramatically under the three-agent setup for each model. GPT-4.1’s privacy preservation rate leaps from 71.2% to 89.7% ( = 18.5 pp), and its adjusted information preservation climbs from 69.9% to 89.5%—an almost 20 pp gain. Across the board, every model records at least a 12 pp increase in privacy preservation, which demonstrates the effectiveness and generalizability of our three-agent approach on the PrivacyLen benchmark.
Appendix E Three-Agent Propagation Analysis

### v2

@@ 5.1 Multi-Agent and Information Flow
To systematically evaluate how different agent configurations balance privacy preservation and completeness of public content, we conducted experiments on single-agent, two-agent, and three-agent pipelines with different information flow configurations. We show our information flow exploration results in Table˜2, with detailed description and case studies are in Appendix˜B. We additionally include a Chain-of-Thought single-agent baseline, which follows the same instructions as the three-agent setup, which detecting events, classifying them as private or public, and generating meeting summaries.
@@
Our results are shown as Table˜2, the multi-agent framework balances secret leakage with public information omission, enhancing both data security and retention. Across both backbones, multi‑agent pipelines substantially reduce our composite error (Leaks Secret or Omits Info↓) relative to single‑agent baselines. For LLaMA-3.1-70B-Instruct, the best configuration is a three‑agent, the information flow that the checker only delivers public information (20.5), improving over the single-agent baseline by 28.0-28.0. For GPT-4o, the best configuration is a three‑agent, Annotate‑Private, no transcript flow ($17.5$), improving over the single‑agent baseline ($26.3$) by 8.8-$8.$ Notably, moving from two to three agents yields additional gains for both models (LLaMA: 24.5 $\rightarrow$ 20.5; GPT‑4o: 21.5 $\rightarrow$ 17.5 depending on flow), indicating that introducing a dedicated checker and clearly separating agent duties improves the privacy utility trade-off provided that each role is well-defined and properly aligned with the task structure.
[…]
LLaMA-3.1-70B-Instruct
single agent vs. three agent (public only) Leakage of private info $\downarrow$ –26.5 pp 11.6 [4.65, 28.92] <0.001<0.001
Omission of public info $\downarrow$ –4.0 pp 2.33 [0.90, 6.07] 0.115
three agent (privacy annotate) vs. three agent (public only) Leakage of private info $\downarrow$ 21.0 pp 11.5 [4.14, 31.95] <0.001<0.001
Omission of public info $\downarrow$ 4.0 pp 2.14 [0.874, 5.256] 0.134
GPT-4o
single agent vs. three agent (public only) Leakage of private info $\downarrow$ –12.0 pp 0.1 [0.02, 0.33] <0.001<0.001
Omission of public info $\downarrow$ 3.0 pp 1.5 [0.72, 3.11] 0.362
three agent (privacy annotate) vs. three agent (public only) Leakage of private info $\downarrow$ –10.0 pp 0.2 [0.06, 0.48] <0.001<0.001
Omission of public info $\downarrow$ +11.0 pp 2.70 [1.42, 5.09] 0.002
[…]
@@
For LLaMA‑3.1‑70B‑Instruct, providing the executor with the meeting transcript leads to a clear advantage for the Public Only strategy: both in the two-agent (Composite error: $24.5$) and three-agent (Composite: $20.5$) settings, this approach outperforms Annotate Private ($34.5$ and $42.0$, respectively). Here, restricting the checker to only forward public information, combined with transcript access, allows the executor to recover useful details and reduces the risk of private information leakage. Importantly, the Public Only strategy also simplifies the executor’s task by eliminating the need for the model to further distinguish between private and public content during its own processing. However, in this setting, if the meeting transcript is withheld from the executor, the performance of Public Only degrades sharply, where omissions of public information increase substantially (e.g., three-agent omission: $19.5$ with transcript vs. $26.0$ without), and composite errors rise (three-agent: 20.5 $\rightarrow$ 32.0). This suggests that LLaMA relies heavily on the transcript to supplement the limited information from the checker, and without it, relevant content is easily lost.
For GPT-4o, the interplay is more balanced. With transcript access, Annotate Private achieves the lowest privacy leakage (5.0), while Public Only achieves the lowest public omission (9.0). When the transcript is withheld, however, the three-agent Annotate Private setting yields the best overall performance (Composite: 17.5), with much lower omission (10.0) than Public Only without transcript (8.5), and still reasonable control of leakage (8.5 vs. 15.0). This indicates that the stronger GPT-4o executor can effectively use checker tags to reconstruct public content, even without seeing the original transcript, and is less prone to privacy leaks than LLaMA in the same setup.
@@
(a) LLaMA-3.1-70B-Instruct
(b) GPT-4o
Figure 2: Comparison of private and public information propagation across agent stages under Public Only and Privacy Annotated settings for both LLaMA-3.1-70B-Instruct and GPT-4o models.
@@ 5.2 Three Agent Information Propagation
To further understand the conditions under which the three-agent framework improves privacy preservation, we quantitatively analyze how information flows across agents and where privacy risks are introduced or mitigated.
@@
Our results (see Figure˜16) demonstrate that the Extractor agent alone is insufficient for reliably identifying public information, necessitating a dedicated Checker agent for effective public information restoration. In this respect, GPT-4o outperforms Llama3.1-70B (55.0% vs. 46.5%) in recovering public content, and the Executor also contributes to the restoration process.
The Checker agent highlights the pronounced superiority of GPT-4o in privacy preservation compared to Llama3.1-70B, as evidenced by a substantially lower privacy leakage rate (33.0% vs. 79.5%). Moreover, GPT-4o achieves a much lower overall information flow leakage rate across the entire pipeline (20.5% vs. 67.0%), which directly substantiates our claim and elucidates why Llama3.1-70B, when used as a base model without providing the meeting transcript to the Executor, suffers significant performance degradation. In contrast, GPT-4o’s performance remains stable,or even improves, under these constraints, illustrating the effectiveness of the Checker’s filtering mechanism and reducing dependence on the downstream agent’s ability to reconstruct prior context (see Figure˜17).
Overall, the composite quality score further reinforces this pattern: while both models exhibit comparable performance at the Extractor stage, Llama3.1-70B’s quality declines substantially at the Checker stage, whereas GPT-4o not only maintains but improves its performance, with a further increase observed at the Executor stage.
@@
To further dissect the role of checker annotation strategies, we conducted a detailed information propagation analysis, comparing Privacy Annotated and Public Only settings across agent stages, and examined how annotation guidance influences privacy preservation and content retention patterns. Statistical significance test results are in Table˜3.
For Llama-3.1-70B-Instruct, our results see(Figure˜2) reveal a striking difference: the proportion of private information included in the Public Only pipeline is substantially higher than in the Privacy Annotated configuration (83.5% vs. 31.0%). This suggests that without explicit privacy annotations, the Public Only setting often fails to sufficiently filter sensitive content. For public information, as expected, the Public Only setting consistently yields higher mention rates across all agent stages compared to the Privacy Annotated setting. In contrast, GPT-4o exhibits an even more pronounced effect (see Figure˜2): when the checker operates in Privacy Annotated mode, the rate of private information mentions is much higher than in Public Only (83% vs. 25.5%). Similarly, for public information, the Public Only setting again achieves higher mention rates than Privacy Annotated.
McNemar tests (see Table˜3) indicate that private-leakage differences between Public Only and Privacy Annotated are consistently statistically reliable across both backbones and comparisons, whereas public-omission effects are weaker and model-dependent: for Llama-3.1-70B-Instruct they are not statistically significant under our matching scheme, while for GPT-4o the Annotated vs. Public Only contrast shows a significant difference in public-content retention.
These results highlight the importance of using a strong model as the checker for effective privacy filtering, while also suggesting that multi-agent pipelines should adaptively assign stronger models to critical validation tasks, with additional verification steps as needed. If the executor is high-capacity and the transcript is hidden, prefer Privacy-Annotated; otherwise prefer Public-Only. Pair either choice with an explicit checker and adversarial leakage tests.
@@ 5.3 Three Agent In Privacylens
Figure˜3 presents the privacy–utility Pareto frontier across all evaluated configurations. We observe a consistent pattern: multi-agent information flows substantially increase privacy preservation with only minor reductions in helpfulness. For both GPT-4.1 and GPT-4o, moving from Single agent to Privacy Annotation or Public Only achieves PPR gains of 15–19 percentage points while maintaining high AHS (from 2.56 to 2.53). Notably, the Privacy Annotation variant offers the best overall Pareto-optimal balance, achieving both high utility and robust privacy protection.
@@ 6 Conclusion
We introduced a multi-agent approach that partitions the tasks of event extraction, classification, and final summary generation among separate agents, addressing the limitations of a single LLM operating alone. Experimental results and ablation studies confirm that our multi-agent pipelines significantly reduce private information leakage without substantially compromising the quality. Notably, we observed performance differences across various information flow configurations and uncovered the multi-agent system’s capacity for self-correction. Our findings highlight the importance of modular, intermediate validation steps in complex, context-dependent scenarios and provide evidence that the choice of checker model (capacity and alignment) and the information flowing materially influence the privacy–utility trade-off, offering practical guidance for multi-agent system design.
@@ 7 Limitation
Our experiments focus on meeting summarization and privacy scenarios defined by the ConfAIde (Mireshghallah et al., 2023) and PrivacyLens (Shao et al., 2025) benchmark. Although our multi‑agent architecture demonstrates strong performance in these controlled benchmarks, domain adaptation remains non‑trivial. The Extractor’s event schema and the Checker’s privacy rules must be handcrafted for each application domain—medical, legal, or financial—each of which has distinct notions of sensitive versus permissible disclosures. This tailoring demands specialized prompt engineering, rule design, and potential collaboration with domain experts. To date, no publicly available benchmarks systematically evaluate multi‑agent or information‑flow architectures for privacy in domains such as healthcare, finance, or legal document processing. Future work should prioritize building such benchmarks, defining domain‑specific privacy norms which would help to better showing the generalizability of our multi-agent system.
@@ A.1 Prompt For ConfAIde
You are an AI assistant that helps summarize conversations from a meeting based on the transcript. The summary will be automatically sent to all the attendees of the meeting, no matter when they joined.
@@ A.3 Hyperparameter
To investigate information propagation, we adopted the following decoding configurations across all experiments for both LLaMA-3-Instruct and GPT-4 models. We set the decoding temperature to 1.0 and used a top-p value of 1.0 to enable unconstrained sampling from the full probability distribution. On ConfAIde, to avoid introducing biases into the generation process, we did not apply frequency or presence penalties (both set to 0.0). For reproducibility, a fixed random seed of 99 was used throughout. Additionally, we generated 10 samples per input prompt to account for generation variability. When evaluating the generalizability of our three-agent framework, we aligned all model decoding temperatures to 1.0 to ensure a fair comparison with the thinking models (i.e., o3 and o4-mini) which use the same sampling configuration.
@@ B.1 Case Studies of the Single Agent Baseline Framework
We present a case study on baseline-agent results in Table˜6. In this scenario, we observe that the single agent fails to identify private information, leading to the private information leakying. (specifically, the private event: a surprise party).
@@ B.2 Case Studies of the Two-Agent Framework
We present a case study on two-agent information flow in Table˜7. In this scenario, we observe that the assistant agent fails to identify private information, leading to the propagation of erroneous data (specifically, the private event: a surprise party) to the subsequent executor agent.
@@ B.3 Case Studies of the Three-Agent Framework
Qualitative case studies, such as the example in Table˜8 and Table˜9, demonstrate the effectiveness of our three-agent configuration. We observe that the Checker Agent, whether filtering events to include only public information or annotating private information, consistently achieves strong performance in identifying and classifying events. This careful segmentation of sensitive versus non-sensitive content results in an Executor Agent output that reliably maintains privacy while conveying essential public details.
@@
Based on the results from the private-only and public-only configurations, we observed that the Executor can autonomously refine the information transmitted by the Checker Agent by leveraging cues from the meeting transcript during the summary generation process. Consequently, we explored a scenario in which the Executor is provided solely with the Checker’s information—omitting the meeting transcript entirely—to assess both the Executor’s performance in this context and the accuracy of the Checker’s public-only setup. Our qualitative investigation and case study (see Table˜10) shows that the Checker often fails to accurately identify all public information, which increases the likelihood of public details being omitted in the final output.
Appendix C Three-Agent Performance Analysis
@@
To probe the single‐agent’s upper bound, in ConfAIde, we introduce a more targeted prompt—Enhanced Single Agent (see Appendix˜A)—designed to maximize public‐information inclusion. Figure˜14 illustrates the trend across four configurations: Baseline Single Agent, Enhanced Single Agent, Three-Agent Privacy Annotation, and Three-Agent Public-Only. Across all models, our Three-Agent framework consistently outperforms the Baseline Single Agent. When compared against the Enhanced Single Agent prompt, the Three-Agent framework delivers especially pronounced gains for open-source models, where it matches or slightly exceeds the performance of the closed-source thinking models. Beyond public inclusion, the Three-Agent framework also shows strong privacy protection across most backbones. Its overall performance is comparable to that of the best closed-source thinking models, while significantly outperforming single-agent settings on non-thinking models.
@@
The privacy-annotated information and public-only information flow’s results of PrivacyLens are shown in Figure˜15. We observe a clear trade‐off between helpfulness and privacy preservation in our three‐agent framework. Our quantitative results show that the binary helpfulness rate and the average helpfulness score remain virtually unchanged when moving from a single‐agent to a three‐agent configuration across all six models. In stark contrast, privacy preservation metrics improve dramatically under the three-agent setup for each model. GPT-4.1’s privacy preservation rate leaps from 71.2% to 89.7% ( = 18.5 pp), and its adjusted information preservation climbs from 69.9% to 89.5%—an almost 20 pp gain. Across the board, every model records at least a 12 pp increase in privacy preservation, which demonstrates the effectiveness and generalizability of our three-agent approach on the PrivacyLen benchmark.
Appendix D Three-Agent Propagation Analysis